A coil of wire contains three loops. If the magnetic flux through the coil changes from −45 Wb to +52 Wb in 0.36 s, what is the magnitude of the emf induced in the coil


1
Expert's answer
2021-05-13T18:02:25-0400

"\phi_1 = -45Wb"


"\phi_2 = 52Wb"


"t = 0.36s"


According to Faraday's law of magnetic induction:


"e = -N\dfrac{d \phi}{dt}"


"e = -3 \times \dfrac{52-(-45)}{0.36-0}"



"e = -3 \times 269.44"


"e = 808.33V"